# Code Equivalence, Point Set Equivalence, and Polynomial Isomorphism

### v1

@@ 7. The Macaulay Inverse System
Let $\mathbb{X}$} be a point set associated to a projective linear $n,k,d]_{q}-$code $C$ with a data rate $k/n$ which is at least $1/2$ and with a reasonably large minimal distance. The minimal distance corresponds to the uniformity of the point set $\mathbb{X}$} (see [28], Prop. 5.11). For sets of points satisfying certain uniformity conditions, bounds for the regularity index $r_{\mathbb{X}}$ are available (see [34] and [33, Sect. 2]).
For instance, the condition that any $k$ points of $\mathbb{X}$} span ${\mathbb{P}}^{k-1}$ is commonly called linearly general position. For points in linearly general position, the corollary after [27, Thm. 5] yields $\Delta\mathop{\rm HF}\nolimits_{\mathbb{X}}(i)\geq k-1$ for $1\leq i\leq r_{\mathbb{X}}-1$. Combining this with our assumption on the data rate, we get $2k\geq n=\sum_{i=0}^{r_{\mathbb{X}}}\Delta\mathop{\rm HF}\nolimits_{\mathbb{X}}(i)\geq(r_{\mathbb{X}}-1)(k-1)+2$, and therefore $r_{\mathbb{X}}\leq 3$. This means that we are applying our formula to polynomials of degree $2r_{\mathbb{X}}-1\leq 5$, and since $\dim_{{\mathbb{F}}_{q}}(P_{5})=\binom{k+4}{5}$, the entire calculation is carried out in polynomial time. In other words, if $\mathbb{X}$} is the associated point set of a “good” linear code with a reasonable data rate and minimal distance, the computation of the Macaulay inverse polynomial can be effected in polynomial time.
Finally, we are ready for the last contribution of this section. A similar result in the local setting was shown in [25, Prop. 2.14].
@@ 8. Point Set Equivalence and Polynomial Isomorphism
Let $C,C^{\prime}$ be two projective $n,k]_{q}-$codes in ${\mathbb{F}}_{q}^{n}$ with generators matrices $G,G^{\prime}\in\mathop{\rm Mat}\nolimits_{k,n}({\mathbb{F}}_{q})$, respectively, let ${\mathbb{X}},{\mathbb{X}}^{\prime}$ be the associated point sets in ${\mathbb{P}}^{k-1}$, and let $\varPhi_{{\mathbb{X}}},\varPhi_{{\mathbb{X}}^{\prime}}\in{\mathbb{F}}_{q}[\pi_{1},\dots,\pi_{k}]$ be their Macaulay inverse polynomials.
@@
Let us prove (a) and (b) simultaneously. Suppose that there exists a linear change of coordinates $\Lambda:\;{\mathbb{P}}^{k-1}\longrightarrow{\mathbb{P}}^{k-1}$ such that $\Lambda({\mathbb{X}})={\mathbb{X}}^{\prime}$. Let $A\in\mathop{\rm GL}\nolimits_{k}({\mathbb{F}}_{q})$ be the matrix defining $\Lambda$, and let $\lambda:\;{\mathbb{F}}_{q}[x_{1},\dots,x_{k}]\longrightarrow{\mathbb{F}}_{q}[x_{1},\dots,x_{q}]$ be the linear change of coordinates defined by $A^{-1}$. By Theorem 6.5, the map $\lambda$ induces an ${\mathbb{F}}_{q}$-algebra isomorphism $\lambda_{D}:\;D_{\mathbb{X}}\longrightarrow D_{{\mathbb{X}}^{\prime}}$, and by Theorem 7.8, the Macaulay inverse polynomials satisfy $\varPhi_{\mathbb{X}}=A^{-1}\ast\varPhi_{{\mathbb{X}}^{\prime}}$ in ${\mathcal{D}}_{-2r_{\mathbb{X}}+1}={\mathbb{F}}_{q}[\pi_{1},\dots,\pi_{k}]_{-2r_{\mathbb{X}}+1}$.
Using Remark 2.14 and Proposition 5.3, we can calculate the vanishing ideals $I_{\mathbb{X}},\,I_{{\mathbb{X}}^{\prime}}$ and system of generators of ${\mathfrak{J}}_{R/{\mathbb{F}}_{q}[\ell]}$ and ${\mathfrak{J}}_{R^{\prime}/{\mathbb{F}}_{q}[\ell^{\prime}]}$ in polynomial time. Thus we get systems of generators of the ideals $\widehat{J}_{\mathbb{X}}$ and $\widehat{J}_{{\mathbb{X}}^{\prime}}$ in $P$ which define $D_{\mathbb{X}}$ and $D_{{\mathbb{X}}^{\prime}}$ in polynomial time.
@@ 9. The Linear Code Equivalence Problem for Iso-Dual Codes
Let ${\mathbb{X}},{\mathbb{X}}^{\prime}$ be two arithmetically Gorenstein set of ${\mathbb{F}}_{q}$-rational points in ${\mathbb{P}}^{k-1}$ such that $\Delta\mathop{\rm HF}\nolimits_{\mathbb{X}}=\Delta\mathop{\rm HF}\nolimits_{{\mathbb{X}}^{\prime}}:\;1\;k{-}1\;k{-}1\;1$. Let ${\varphi}_{\mathbb{X}},{\varphi}_{{\mathbb{X}}^{\prime}}\in{\mathcal{D}}_{-3}$ be the homogeneous elements generating the Macaulay inverse systems of $\overline{R}$ and $\overline{R}^{\prime}$, respectively.
@@
This proposition yields the following polynomial time reduction.
@@
In the proof of the proposition, the ideals $I_{\mathbb{X}}+\langle L\rangle$ and $I_{{\mathbb{X}}^{\prime}}+\langle L^{\prime}\rangle$ defining $R$ and $R^{\prime}$, respectively, can be found in polynomial time by Remark 2.14. Then vector space bases of their homogeneous components of degree 3 can be found in polynomial time, as $\dim_{{\mathbb{F}}_{q}}(P_{3})=\binom{k+2}{3}$. Finally, the calculation of the dual vector space in ${\mathbb{F}}_{q}[\pi_{1},\dots,\pi_{k}]_{3}$ with respect to the non-degenerate bilinear map $\circ_{3}$ in Proposition 7.1.d proceeds using linear algebra in polynomial time. ∎

### v2

@@ 7. The Macaulay Inverse System
Let $K={\mathbb{F}}_{q}$, and let $\mathbb{X}$} be a point set associated to a projective linear $n,k,d]_{q}-$code $C$ with a data rate $k/n$ which is at least $1/2$ and with a reasonably large minimal distance. The minimal distance corresponds to the uniformity of the point set $\mathbb{X}$} (see [36], Prop. 5.11). For sets of points satisfying certain uniformity conditions, bounds for the regularity index $r_{\mathbb{X}}$ are available (see [42] and [41, Sect. 2]).
For instance, the condition that any $k$ points of $\mathbb{X}$} span ${\mathbb{P}}^{k-1}$ is commonly called linearly general position. For points in linearly general position, the corollary after [35, Thm. 5] yields $\Delta\mathop{\rm HF}\nolimits_{\mathbb{X}}(i)\geq k-1$ for $1\leq i\leq r_{\mathbb{X}}-1$. Combining this with our assumption on the data rate, we get $2k\geq n=\sum_{i=0}^{r_{\mathbb{X}}}\Delta\mathop{\rm HF}\nolimits_{\mathbb{X}}(i)\geq(r_{\mathbb{X}}-1)(k-1)+2$, and therefore $r_{\mathbb{X}}\leq 3$. This means that we are applying our formula to polynomials of degree $2r_{\mathbb{X}}-1\leq 5$, and since $\dim_{{\mathbb{F}}_{q}}(P_{5})=\binom{k+4}{5}$, the entire calculation is carried out in polynomial time. In other words, if $\mathbb{X}$} is the associated point set of a “good” linear code with a reasonable data rate and minimal distance, the computation of the Macaulay inverse polynomial can be effected in polynomial time.
Finally, we are ready for the last contribution of this section. A similar result in the local setting was shown in [33, Prop. 2.14]. Here we let $K={\mathbb{F}}_{q}$ again.
@@ 8. Point Set Equivalence and Polynomial Isomorphism
Let $C,C^{\prime}$ be two projective $n,k]_{q}-$codes in ${\mathbb{F}}_{q}^{n}$ with generator matrices $G,G^{\prime}\in\mathop{\rm Mat}\nolimits_{k,n}({\mathbb{F}}_{q})$, respectively, let ${\mathbb{X}},{\mathbb{X}}^{\prime}$ be the associated point sets in ${\mathbb{P}}^{k-1}$, and let $\varPhi_{{\mathbb{X}}},\varPhi_{{\mathbb{X}}^{\prime}}\in{\mathbb{F}}_{q}[\pi_{1},\dots,\pi_{k}]$ be their Macaulay inverse polynomials.
@@
Let us prove (a) and (b) simultaneously. Suppose that there exists a linear change of coordinates $\Lambda:\;{\mathbb{P}}^{k-1}\longrightarrow{\mathbb{P}}^{k-1}$ such that $\Lambda({\mathbb{X}})={\mathbb{X}}^{\prime}$. Let $A\in\mathop{\rm GL}\nolimits_{k}({\mathbb{F}}_{q})$ be the matrix defining $\Lambda$, and let $\lambda:\;{\mathbb{F}}_{q}[x_{1},\dots,x_{k}]\longrightarrow{\mathbb{F}}_{q}[x_{1},\dots,x_{k}]$ be the linear change of coordinates defined by $A^{-1}$. By Theorem 6.5, the map $\lambda$ induces an ${\mathbb{F}}_{q}$-algebra isomorphism $\lambda_{D}:\;D_{\mathbb{X}}\longrightarrow D_{{\mathbb{X}}^{\prime}}$, and by Theorem 7.8, the Macaulay inverse polynomials satisfy $\varPhi_{\mathbb{X}}=A^{-1}\ast\varPhi_{{\mathbb{X}}^{\prime}}$ in ${\mathcal{D}}_{-2r_{\mathbb{X}}+1}={\mathbb{F}}_{q}[\pi_{1},\dots,\pi_{k}]_{-2r_{\mathbb{X}}+1}$.
Using Remark 2.14 and Proposition 5.3, we can calculate the vanishing ideals $I_{\mathbb{X}},\,I_{{\mathbb{X}}^{\prime}}$ and systems of generators of ${\mathfrak{J}}_{R/{\mathbb{F}}_{q}[\ell]}$ and ${\mathfrak{J}}_{R^{\prime}/{\mathbb{F}}_{q}[\ell^{\prime}]}$ in polynomial time. Thus we get systems of generators of the ideals $\widehat{J}_{\mathbb{X}}$ and $\widehat{J}_{{\mathbb{X}}^{\prime}}$ in $P$ which define $D_{\mathbb{X}}$ and $D_{{\mathbb{X}}^{\prime}}$ in polynomial time.
@@ 9. The Linear Code Equivalence Problem for Iso-Dual Codes
Let ${\mathbb{X}},{\mathbb{X}}^{\prime}$ be two arithmetically Gorenstein sets of ${\mathbb{F}}_{q}$-rational points in ${\mathbb{P}}^{k-1}$ such that $\Delta\mathop{\rm HF}\nolimits_{\mathbb{X}}=\Delta\mathop{\rm HF}\nolimits_{{\mathbb{X}}^{\prime}}:\;1\;k{-}1\;k{-}1\;1$. Let ${\varphi}_{\mathbb{X}},{\varphi}_{{\mathbb{X}}^{\prime}}\in{\mathcal{D}}_{-3}$ be the homogeneous elements generating the Macaulay inverse systems of $\overline{R}$ and $\overline{R}^{\prime}$, respectively.
@@
Together with the following result, this proposition yields a polynomial time reduction from the PSE seach problem to the 3-PI search problem, as explained below. The explicit computation of the Macaulay inverse polynomial of the Artinian reduction of an arithmetically Gorenstein set of points has been described before in [24], and when the base field has characteristic zero in more detail in [47] and [23]. The next proposition may be seen as a slightly improved characteristic free version of these results.
Proposition 9.2 (Computing the Macaulay Inverse Polynomial of an Artinian Reduction of an Arithmetically Gorenstein Set of Points).
Let $K$ be a field, let ${\mathbb{X}}=\{p_{1},\dots,p_{n}\}$ be a set of $K$-rational points in ${\mathbb{P}}^{k-1}$, and assume that $x_{1}\in R_{1}$ is a non-zerodivisor. Let $\overline{R}=R/\langle x_{1}\rangle$ and write $p_{i}=(1:p_{i2}:\cdots:p_{ik})$ with $p_{ij}\in K$ for $i=1,\dots,n$. For $i=1,\dots,n$, let $L_{i}=\pi_{1}+p_{i2}\pi_{2}+\cdots+p_{ik}\pi_{k}\in{\mathcal{D}}_{-1}$.
Assume that $\mathbb{X}$} is arithmetically Gorenstein. Let $f_{1},\dots,f_{n}\in R_{r_{\mathbb{X}}}$ be the separators of $\mathbb{X}$} and write $\bar{f}_{j}=\beta_{j}\,\bar{f}_{1}$ with $\beta_{j}\in K\setminus\{0\}$ for $j=2,\dots,k$, where $\bar{f}_{i}$ is the residue class of $f_{i}$ in $\overline{R}_{r_{\mathbb{X}}}$.
In this setting, a Macaulay inverse polynomial ${\varphi}_{\mathbb{X}}$ of $\mathbb{X}$} is given by
${\varphi}_{\mathbb{X}}\;=\;L_{1}^{[r_{\mathbb{X}}]}+\beta_{2}L_{2}^{[r_{\mathbb{X}}]}+\cdots+\beta_{n}L_{n}^{[r_{\mathbb{X}}]}\in{\mathcal{D}}_{-r_{\mathbb{X}}}$
Here we let $L_{i}^{[j]}=\sum_{|\alpha|=j}p_{i}^{\alpha}\pi^{\alpha}$ for all $i=1,\dots,n$ and $j\geq 0$.
Proof.
By [24, Thm. I], there exists a tuple $(\alpha_{1},\dots,\alpha_{n})\in K^{n}$ such that ${\varphi}_{\mathbb{X}}=\alpha_{1}L_{1}^{[r_{\mathbb{X}}]}+\cdots+\alpha_{n}L_{n}^{[r_{\mathbb{X}}]}$. Now we use the formula $h\circ L_{i}^{[j]}=L_{i}^{[j-r]}\cdot h(p_{i})$ for $h\in R_{r}$ given in [24, Sec. 2, Lemma]. It shows that $f_{i}\circ L_{j}^{[r_{\mathbb{X}}]}=\delta_{ij}$ for all $i,j$, and therefore $f_{i}\circ{\varphi}_{\mathbb{X}}=\alpha_{i}$ for $i=1,\dots,n$.
As shown in [47, Prop. 2.2], we have $x_{1}\circ{\varphi}_{\mathbb{X}}=\alpha_{1}L_{1}^{[r_{\mathbb{X}}-1]}+\cdots+\alpha_{n}L_{n}^{[r_{\mathbb{X}}-1]}=0$. (The assumption ${\rm char}(K)=0$ made there plays no role in the proof.) Consequently, if we write $f_{j}-\beta_{j}f_{1}=x_{1}g_{j}$ with $g_{j}\in R_{r_{\mathbb{X}}-1}$ for $j=2,\dots,n$, then
$0\;=\;g_{j}\circ(x_{1}\circ{\varphi}_{\mathbb{X}})\;=\;(x_{1}g_{j})\circ{\varphi}_{\mathbb{X}}=(f_{j}-\beta_{j}f_{1})\circ{\varphi}_{\mathbb{X}}\;=\;(\alpha_{j}-\beta_{j}\alpha_{1})\circ{\varphi}_{\mathbb{X}}$
shows $\alpha_{j}=\beta_{j}\alpha_{1}$ for $j=2,\dots,n$. As the tuple $(\alpha_{1},\dots,\alpha_{n})$ is determined only up to a non-zero scalar factor, this proves the claim. ∎
Given a bound for $r_{\mathbb{X}}$, the polynomial ${\varphi}_{\mathbb{X}}$ can be calculated efficiently. This results in the following polynomial time reduction.
@@
In the proof of Prop. 9.1, the ideals $I_{\mathbb{X}}+\langle L\rangle$ and $I_{{\mathbb{X}}^{\prime}}+\langle L^{\prime}\rangle$ defining $R$ and $R^{\prime}$, respectively, can be found in polynomial time by Remark 2.14. Then vector space bases of their homogeneous components of degree 3 can be found in polynomial time, as $\dim_{{\mathbb{F}}_{q}}(P_{3})=\binom{k+2}{3}$. Finally, the calculation of the Macaulay dual polynomials ${\varphi}_{\mathbb{X}},{\varphi}_{{\mathbb{X}}^{\prime}}\in{\mathbb{F}}_{q}[\pi_{1},\dots,\pi_{k}]_{3}$ is achieved in polynomial time using the preceding proposition. ∎